# A Survey on 3D Gaussian Splatting

### v8

@@ 3.2.1. Parameter Optimization
where $\lambda\in[0,1]$ is a weighting factor.
@@
where $\bm{R}$ is the rotation matrix derived from the quaternion $\bm{q}$, and $\bm{S}$ is the scaling matrix given by $\text{diag}(\bm{s})$. As seen, there is a complex computational graph to obtain the opacity $\alpha$, i.e., $\bm{q}$ and $\bm{s}\mapsto\bm{\Sigma}$, $\bm{\Sigma}\mapsto\bm{\Sigma}^{\prime}$, and $\bm{\Sigma}^{\prime}\mapsto\alpha$. To avoid the cost of automatic differentiation, 3D GS derives the gradients for $\bm{q}$ and $\bm{s}$ so as to compute them directly during optimization.
@@ 3.2.2. Density Control
Though 3D GS has achieved impressive milestones, significant room for improvement remains, e.g., data and hardware requirement, rendering and optimization algorithm, and applications in downstream tasks. In the subsequent sections, we seek to elaborate on select extended versions. These are: i) 3D GS for Sparse Input (Xiong et al., 2023; Zhu et al., 2024a; Charatan et al., 2024; Szymanowicz et al., 2024b; Li et al., 2024i; Swann et al., 2024; Chen et al., 2024e; Wewer et al., 2024; Xu et al., 2024b; Shen et al., 2024; Zhang et al., 2024f) (Sec. 4.1), ii) Memory-efficient 3D GS (Fan et al., 2023; Navaneet et al., 2023; Lee et al., 2024c; Morgenstern et al., 2023; Zhang et al., 2024d; Niedermayr et al., 2024; Chen et al., 2024c; Papantonakis et al., 2024; Fang and Wang, 2024) (Sec. 4.2), iii) Photorealistic 3D GS (Yu et al., 2024a; Gao et al., 2023; Yan et al., 2024b; Jiang et al., 2024b; Lee et al., 2024b; Malarz et al., 2023; Bolanos et al., 2024; Radl et al., 2024; Yang et al., 2024a; Peng et al., 2024b; Zhao et al., 2024b; Dahmani et al., 2024; Li et al., 2024f; Liang et al., 2024b; Seiskari et al., 2024; Song et al., 2024) (Sec. 4.3), iv) Improved Optimization Algorithms (Fu et al., 2024; Lu et al., 2024b; Jung et al., 2024; Li et al., 2024f; Yu et al., 2024c; Zhang et al., 2024i; Huang et al., 2024a; Li et al., 2024c) (Sec. 4.4), v) 3D Gaussian with More Properties (Shi et al., 2024; Qin et al., 2024; Zuo et al., 2024; Zhou et al., 2024a; Ye et al., 2024; Cen et al., 2023; Yang et al., 2024d) (Sec. 4.5), vi) Hybrid Representation (Yang et al., 2024b; Wu et al., 2024c; Xu et al., 2024a) (Sec. 4.6), and vii) New Rendering Algorithm (Sec. 4.7). While we have carefully selected several key directions, we acknowledge that it is inevitably a biased view. A more comprehensive collection is given in Github.
@@
Existing methods can be categorized into two primary groups. i) Regularization based methods introduce additional constraints such as depth information to enhance the detail and global consistency (Li et al., 2024i; Zhu et al., 2024a; Chen et al., 2024e; Zhang et al., 2024f). For example, DNGaussian (Li et al., 2024i) introduced a depth-regularized approach to address the challenge of geometry degradation in sparse input. FSGS (Zhu et al., 2024a) devised a Gaussian Unpooling process for initialization and also introduced depth regularization. MVSplat (Chen et al., 2024e) proposed a cost volume representation so as to provide geometry cues. Unfortunately, when dealing with a limited number of views, or even just one, the efficacy of regularization techniques tends to diminish, which leads to ii) generalizability based methods that use learned priors (Charatan et al., 2024; Szymanowicz et al., 2024b; Xu et al., 2024b; Szymanowicz et al., 2024a). One approach involves synthesizing additional views through generative models, which can be seamlessly integrated into existing reconstruction pipelines (Sargent et al., 2024). However, this augmentation strategy is computationally intensive and inherently bounded by the capabilities of the used generative model. Another well-known paradigm employs feed-forward Gaussian model to directly generates the properties of a set of 3D Gaussians. This paradigm typically requires multiple views for training but can reconstruct 3D scenes with only one input image. For instance, PixelSplat (Charatan et al., 2024) proposed to sample Gaussians from dense probability distributions. Splatter Image (Szymanowicz et al., 2024b) introduced a 2D image-to-image network that maps an input image to a 3D Gaussian per pixel. However, as the generated pixel-aligned Gaussians are distributed nearly evenly in the space, they struggle to represent high-frequency details and smoother regions with an appropriate number of Gaussians.
The challenge of 3D GS for sparse inputs centers on the modeling of priors, whether through depth information, generative models, or feed-forward Gaussian models. The fundamental trade-off lies between overfitting to available views and using learned priors for generalization. Future research could explore adaptive mechanisms for controlling this trade-off, potentially through learned confidence measures, context-aware prior selection, user preferences, etc. In addition, while current methods focus on static scenes, extending these approaches to dynamic scenarios presents an exciting frontier for investigation, particularly in handling temporal consistency and motion-induced artifacts.
@@ 4.2. Memory-efficient 3D GS
While 3D GS demonstrates remarkable capabilities, its scalability poses significant challenges, particularly when juxtaposed with NeRF-based methods. The latter benefits from the simplicity of storing merely the parameters of a learned MLP. This scalability issue becomes increasingly acute in the context of large-scale scene management, where the computational and memory demands escalate substantially. Consequently, there is an urgent need to optimize memory usage in both model training and storage.
Recent research has pursued two primary directions to address memory efficiency. First, several approaches focus on reducing the number of 3D Gaussians (Papantonakis et al., 2024; Lee et al., 2024c; Chen et al., 2024c). These methods either employ strategic pruning of low-impact Gaussians, such as the volume-based masking (Lee et al., 2024c), or represent neighboring Gaussians using the same properties stored within a “local anchor” obtained by clustering (Lu et al., 2024b), hash-grid (Chen et al., 2024c), etc. Second, researchers have developed methods for compressing Gaussian’s properties (Niedermayr et al., 2024; Lee et al., 2024c; Chen et al., 2024c). For instance, Niedermayr et al. (Niedermayr et al., 2024) compressed color and Gaussian parameters into compact codebooks, using sensitivity measures for effective quantization and fine-tuning. HAC (Chen et al., 2024c) predicted the probability of each quantized attribute using Gaussian distributions and then devise an adaptive quantization module. These directions are not mutually exclusive; instead, one framework might use a hybrid approach combining multiple strategies.
@@ 4.3. Photorealistic 3D GS
The current rendering pipeline of 3D GS (Sec. 3.1) is straightforward and involves several drawbacks. For instance, the simple visibility algorithm may lead to a drastic switch in the depth/blending order of Gaussians (Kerbl et al., 2023). The visual fidelity of rendered images, including aspects such as aliasing, reflections, and artifacts, can be further optimized.
@@
Recent research has focused on addressing three main aspects of visual quality, with aliasing being specific to 3D GS’s rendering algorithm, while reflection and blur handling represent broader challenges in 3D reconstruction. i) Aliasing. Due to the discrete sampling paradigm (viewing each pixel as a single point instead of an area), 3D GS is susceptible to aliasing when dealing with varying resolutions, which leads to blurring or jagged edges. Solutions emerged at both training and inference stages. Researchers developed training-time improvements from the sampling rate perspective and introduced schemes such as multi-scale Gaussians (Yan et al., 2024b), 2D Mip filter (Yu et al., 2024a), and conditioned logistic function (Liang et al., 2024b). Inference-time solutions, such as 2D scale-adaptive filtering (Song et al., 2024), offer enhanced fidelity that can be integrated into any existing 3D GS frameworks. ii) Reflection. Achieving realistic rendering of reflective materials is a hard, long-standing problem in 3D scene reconstruction. Recent works have introduced various approaches to model reflective materials (Jiang et al., 2024b; Meng et al., 2024; Yang et al., 2024a) and enable relightable Gaussian representation (Saito et al., 2024), though achieving physically accurate specular effects remains challenging. iii) Blur. While 3D GS excels on carefully curated datasets, real-world captures often suffer from blurs such as motion blur and defocus blur. Recent approaches explicitly incorporated blur modeling during training, employing techniques such as coarse-to-fine kernel optimization (Peng et al., 2024b) and photometric bundle adjustment (Zhao et al., 2024b) to address this challenge.
While the approximations made in 3D GS (Sec. 3.1) contribute to its computational efficiency, they also lead to aliasing, difficulties in illumination estimation, etc. Current solutions, though impressive, typically address individual problems rather than providing a universal solution. A practical intermediate approach involves first detecting specific issues (e.g., aliasing, blur) and then applying targeted optimization strategies. The ultimate goal remains developing an advanced reconstruction system that overcomes these limitations, either through fundamental improvements to 3D GS or through brand-new architectures.
@@ 4.4. Improved Optimization Algorithms
Three main directions stand out for improving the optimization of 3D GS. i) Additional Regularization (e.g., frequency (Zhang et al., 2024i) and geometry (Lu et al., 2024b; Li et al., 2024f)). Geometry-aware approaches have been particularly successful, preserving scene structure through the incorporation of local anchor points (Lu et al., 2024b), depth and surface constraints (Chen et al., 2023a; Yu et al., 2024d; Zhang et al., 2024c), Gaussian volumes (Chen et al., 2024d), etc. ii) Optimization Procedure Enhancement (Cheng et al., 2024; Yu et al., 2024d; Ververas et al., 2024). While the original strategy of density control (Sec. 3.2.2) has proven valuable, considerable room for improvement remains. For example, GaussianPro (Cheng et al., 2024) addresses the challenge of dense initialization in texture-less surfaces and large-scale scenes through an advanced Gaussian densification strategy. iii) Constraint Relaxation. Reliance on external tools/algorithms can introduce errors and cap the system’s performance potential. For instance, SfM, commonly used in the initialization process, is error-prone and struggle with complex scenes. Recent works have begun exploring COLMAP-free approaches utilizing stream continuity (Fu et al., 2024; Smith et al., 2024), potentially enabling learning from internet-scale unposed video datasets.
@@
Despite impressive, the properties of 3D Gaussian (Sec. 3.1) are designed to be used for novel-view synthesis only. By augmenting 3D Gaussian with additional properties, such as linguistic (Shi et al., 2024; Qin et al., 2024; Zuo et al., 2024), semantic/instance (Zhou et al., 2024a; Ye et al., 2024; Cen et al., 2023), and spatial-temporal (Yang et al., 2024d) properties, 3D GS demonstrates its considerable potential to revolutionize various domains.
Here we list several interesting applications using 3D Gaussians with specially designed properties. i) Language Embedded Scene Representation (Shi et al., 2024; Qin et al., 2024; Zuo et al., 2024). Due to the high computational and memory demands of current language-embedded scene representations, Shi et al. (Shi et al., 2024) proposed a quantization scheme that augments 3D Gaussian with streamlined language embeddings instead of the original high-dimensional embeddings. This method also mitigated semantic ambiguity and enhanced the precision of open-vocabulary querying by smoothing out semantic features across different views, guided by uncertainty values. ii) Scene Understanding and Editing (Zhou et al., 2024a; Ye et al., 2024; Cen et al., 2023). Feature 3DGS (Zhou et al., 2024a) integrated 3D GS with feature field distillation from 2D foundation models. By learning a lower-dimensional feature field and applying a lightweight convolutional decoder for upsampling, Feature 3DGS achieved faster training and rendering speeds while enabling high-quality feature field distillation, supporting applications like semantic segmentation and language-guided editing. iii) Spatiotemporal Modeling (Yang et al., 2024d; Lin et al., 2024a). To capture the complex spatial and temporal dynamics of 3D scenes, Yang et al. (Yang et al., 2024d) conceptualized spacetime as a unified entity and approximates the spatiotemporal volume of dynamic scenes using a collection of 4D Gaussians. The proposed 4D Gaussian representation and corresponding rendering pipeline are capable of modeling arbitrary rotations in space and time and allow for end-to-end training.
@@ 4.6. Hybrid Representation
Next we showcase various fascinating uses of 3D GS with specially devised structured information. i) Facial Expression Modeling. Considering the challenge of creating high-fidelity 3D head avatars under sparse view conditions, Gaussian Head Avatar (Xu et al., 2024a) introduced controllable 3D Gaussians and an MLP-based deformation field. Concretely, it captured detailed facial expressions and dynamics by optimizing neutral 3D Gaussians alongside the deformation field, thus ensuring both detail fidelity and expression accuracy. ii) Spatiotemporal Modeling. Yang et al. (Yang et al., 2024b) proposed to reconstruct dynamic scenes with deformable 3D Gaussians. The deformable 3D Gaussians are learned in a canonical space, coupled with a deformation field (i.e., a spatial MLP) that models the spatial-temporal dynamics. The proposed method also incorporated an annealing smoothing training mechanism to enhance temporal smoothness without additional computational costs. iii) Style Transfer. Saroha et al. (Saroha et al., 2024) proposed GS in style, an advanced approach for real-time neural scene stylization. To maintain a cohesive stylized appearance across multiple views without compromising on rendering speed, they used pre-trained 3D Gaussians coupled with a multi-resolution hash grid and a small MLP to produce stylized views. In a nutshell, incorporating structured information can serve as a complementary part for adapting to tasks that are incompatible with the sparsity and disorder of 3D Gaussians.

### v9

@@ 3.2.1. Parameter Optimization
where $\lambda\in[0,1]$ is a weighting factor. Here, $\mathcal{L}_{\text{D-SSIM}}$ is formulated as $1-\text{SSIM}$, derived from the structural similarity index measure (SSIM). While $\mathcal{L}_{1}$ minimizes absolute pixel-wise differences, $\mathcal{L}_{\text{D-SSIM}}$ acts as a perceptual metric that enforces structural consistency, ensuring that the synthesized images preserve high-frequency details and local luminance patterns often missed by $\ell_{1}$ alone.
@@
where $\bm{R}$ is the rotation matrix derived from the quaternion $\bm{q}$, and $\bm{S}$ is the scaling matrix given by $\text{diag}(\bm{s})$. As seen, there is a complex computational graph to obtain the opacity $\alpha$, i.e., $\bm{q}$ and $\bm{s}\mapsto\bm{\Sigma}$, $\bm{\Sigma}\mapsto\bm{\Sigma}^{\prime}$, and $\bm{\Sigma}^{\prime}\mapsto\alpha$. To avoid the cost of automatic differentiation, 3D GS derives the gradients for $\bm{q}$ and $\bm{s}$ to compute them directly during optimization.
@@ 3.2.2. Density Control
Though 3D GS has achieved impressive milestones, significant room for improvement remains, e.g., data and hardware requirement, rendering and optimization algorithm, and applications in downstream tasks. In the subsequent sections, we seek to elaborate on select extended versions. These are: i) 3D GS for Sparse Input (Sec. 4.1), ii) Memory-efficient 3D GS (Sec. 4.2), iii) Photorealistic 3D GS (Sec. 4.3), iv) Improved Optimization Algorithms (Sec. 4.4), v) 3D Gaussian with More Properties (Sec. 4.5), vi) Hybrid Representation (Sec. 4.6), and vii) New Rendering Algorithm (Sec. 4.7). To provide a synthesized overview before delving into each direction, Table S1 summarizes their key ideas, representative works (ordered chronologically), advantages, and limitations. While we have carefully selected several key directions, we acknowledge that it is inevitably a biased view. A more comprehensive collection is given in Github.
@@
Existing methods can be categorized into two groups (cf. Table S2). i) Regularization based methods introduce additional constraints such as depth information to enhance the detail and global consistency (Li et al., 2024c; Zhu et al., 2024c; Chen et al., 2024d; Zhang et al., 2024e). For example, DNGaussian (Li et al., 2024c) introduced a depth-regularized scheme to address the challenge of geometry degradation. FSGS (Zhu et al., 2024c) devised a Gaussian Unpooling process for initialization and also introduced depth regularization. MVSplat (Chen et al., 2024d) proposed a cost volume representation so as to provide geometry cues. Unfortunately, when dealing with a limited number of views, or even just one, the efficacy of regularization techniques tends to diminish, which leads to ii) generalizability based methods that use learned priors (Charatan et al., 2024; Szymanowicz et al., 2024b; Xu et al., 2024b; Szymanowicz et al., 2024a). One approach involves synthesizing additional views through generative models, which can be seamlessly integrated into existing reconstruction pipelines (Sargent et al., 2024). However, this augmentation strategy is computationally intensive and inherently bounded by the capabilities of the used generative model. Another well-known paradigm employs feed-forward Gaussian model to directly generates the properties of a set of 3D Gaussians. This paradigm typically requires multiple views for training but can reconstruct 3D scenes with only one input image. For instance, PixelSplat (Charatan et al., 2024) proposed to sample Gaussians from dense probability distributions. Splatter Image (Szymanowicz et al., 2024b) introduced a 2D image-to-image network that maps an input image to a 3D Gaussian per pixel. However, as the generated pixel-aligned Gaussians are distributed nearly evenly in the space, they struggle to represent high-frequency details and smoother regions with an appropriate number of Gaussians.
The challenge of 3DGS for sparse inputs centers on how priors are modeled, whether through depth information, generative models, or feed-forward Gaussian models. In practice, there is a trade-off between regularization-based methods, which rely on strong constraints and per-scene optimization (favoring geometric quality but incurring higher computation), and generalizability-based methods, which use learned priors for fast feed-forward inference (favoring speed and cross-scene generalization, but sometimes at the expense of geometric fidelity). We summarize these characteristics in Table S2. Future research could explore adaptive mechanisms to control this trade-off, for example via learned confidence measures, context-aware prior selection, or user preferences, etc. In addition, while most current methods focus on static scenes, extending these approaches to dynamic scenarios presents an exciting direction, particularly in handling temporal consistency and motion-induced artifacts.
@@ 4.2. Memory-efficient 3D GS
While 3D GS demonstrates remarkable capabilities, its scalability poses significant challenges, particularly when juxtaposed with NeRF-based methods. The latter benefits from the simplicity of storing merely the parameters of a learned MLP. This scalability issue becomes acute in the context of large-scale scene management, where the computational and memory demands escalate substantially. Consequently, there is an urgent need to optimize memory usage in both model training and storage. Please see the excellent surveys (Bagdasarian et al., 2025; Ali et al., 2025) for more insights on GS compression.
Recent research has pursued two primary directions to address memory efficiency. First, several approaches focus on reducing the number of 3D Gaussians (Papantonakis et al., 2024; Lee et al., 2024b; Chen et al., 2024b). These methods either employ strategic pruning of low-impact Gaussians, such as the volume-based masking (Lee et al., 2024b), or represent neighboring Gaussians using the same properties stored within a “local anchor” obtained by clustering (Lu et al., 2024b), hash-grid (Chen et al., 2024b), etc. Second, researchers have developed methods for compressing Gaussian’s properties (Niedermayr et al., 2024; Lee et al., 2024b; Chen et al., 2024b). For instance, Niedermayr et al. (Niedermayr et al., 2024) compressed color and Gaussian parameters into compact codebooks, using sensitivity measures for effective quantization and fine-tuning. HAC (Chen et al., 2024b) predicted the probability of each quantized attribute using Gaussian distributions and then devise an adaptive quantization module. These directions are not mutually exclusive; instead, one framework might use a hybrid approach combining multiple strategies.
@@ 4.3. Photorealistic 3D GS
The current rendering pipeline of 3D GS (Sec. 3.1) is straightforward and involves several drawbacks. For instance, the simple visibility algorithm may lead to a drastic switch in the depth/blending order of Gaussians. The visual fidelity of rendered images, including aspects such as aliasing, reflections, and artifacts, can be further optimized.
@@
Recent research has focused on addressing three main aspects of visual quality, with aliasing being specific to 3D GS’s rendering algorithm, while reflection and blur handling represent broader challenges in 3D reconstruction. i) Aliasing. Due to the discrete sampling paradigm (viewing each pixel as a single point instead of an area), 3D GS is susceptible to aliasing when dealing with varying resolutions, which leads to blurring or jagged edges. Solutions emerged at both training and inference stages. Researchers developed training-time improvements from the sampling rate perspective and introduced schemes such as multi-scale Gaussians (Yan et al., 2024c), 2D Mip filter (Yu et al., 2024c) (predecessor of Mip-NeRF (Barron et al., 2021)), and conditioned logistic function (Liang et al., 2024b). Inference-time solutions, such as 2D scale-adaptive filtering (Song et al., 2024), offer enhanced fidelity that can be integrated into any existing 3D GS frameworks. ii) Challenging objects. Realistic reconstruction of objects with complex appearance, such as reflective, transparent, highly glossy, or anisotropic surfaces, remains a long-standing problem. 3D GS inherits these difficulties because its rasterization-based pipeline only approximates secondary effects. Recent works therefore extend 3D GS with shading functions defined on per-Gaussian normals to better handle reflective and specular surfaces (Jiang et al., 2024b), with mirror attributes and plane-mirror imaging to reconstruct mirror geometry and reflections (Meng et al., 2024), and with anisotropic view-dependent appearance fields that more faithfully capture specular and anisotropic components (Yang et al., 2024c), as well as relightable Gaussian representation (Saito et al., 2024). Despite these advances, robust reconstruction of reflective and transparent objects in general scenes remains challenging. iii) Blur. While 3D GS excels on carefully curated datasets, real-world captures often suffer from blurs such as motion blur and defocus blur. Recent approaches explicitly incorporated blur modeling during training, employing techniques such as coarse-to-fine kernel optimization (Peng et al., 2024a) and photometric bundle adjustment (Zhao et al., 2024c) to address this challenge.
While the approximations made in GS (Sec. 3.1) contribute to its efficiency, they also underlie artifacts such as aliasing and blur, and limit the accuracy of illumination estimation and the modeling of challenging objects. Most existing efforts focus on specific failure modes (e.g., anti-aliasing, deblurring, or handling particular challenging objects) rather than providing a universal solution. In practice, one can first diagnose which issues are present in a scene (e.g., aliasing, blur, or reflective and transparent objects) and then apply targeted optimization strategies. For static object reconstruction in particular, faithfully modeling challenging objects such as reflective, transparent, or highly view-dependent surfaces remains difficult despite recent progress, and extending these capabilities to dynamic or relightable scenes further increases the complexity. The ultimate goal is to develop more general reconstruction systems that alleviate these limitations, either through fundamental improvements to 3D GS or through brand-new architectures.
@@ 4.4. Improved Optimization Algorithms
Three main directions stand out for improving the optimization of 3D GS. i) Additional Regularization (e.g., frequency (Zhang et al., 2024d) and geometry (Lu et al., 2024b; Li et al., 2024h)). Geometry-aware approaches have been particularly successful, preserving scene structure through the incorporation of local anchor points (Lu et al., 2024b), depth and surface constraints (Chen et al., 2023a; Yu et al., 2024d; Zhang et al., 2024a), Gaussian volumes (Chen et al., 2024a), etc. ii) Optimization Procedure Enhancement (Cheng et al., 2024; Yu et al., 2024d; Ververas et al., 2024). While the original strategy of density control (Sec. 3.2.2) has proven valuable, considerable room for improvement remains. For example, GaussianPro (Cheng et al., 2024) addresses the challenge of dense initialization in texture-less surfaces and large-scale scenes through an advanced Gaussian densification strategy. iii) Constraint Relaxation. Reliance on external tools/algorithms can introduce errors and cap the system’s performance potential. For instance, SfM, commonly used in the initialization process, is error-prone and struggle with complex scenes. Recent works have begun exploring COLMAP-free (Schonberger and Frahm, 2016) approaches utilizing stream continuity (Fu et al., 2024; Smith et al., 2024), potentially enabling learning from internet-scale unposed video datasets. This line of work removes the dependency on pre-computed SfM by jointly optimizing camera poses and Gaussian parameters in an offline/batch manner. Readers interested in the online counterpart of this unknown-pose problem (tracking + mapping under real-time constraints) are referred to Sec. 5.1 on GS-based SLAM.
@@
Despite impressive, the properties of 3D Gaussian (Sec. 3.1) are designed to be used for novel-view synthesis only. By augmenting 3D Gaussian with additional properties, such as linguistic (Shi et al., 2024; Qin et al., 2024; Zuo et al., 2024), semantic/instance (Zhou et al., 2024b; Ye et al., 2024; Cen et al., 2023), and spatial-temporal (Yang et al., 2024b) properties, 3D GS demonstrates its considerable potential to revolutionize various domains.
Here we list several interesting applications using 3D Gaussians with specially designed properties. i) Language Embedded Scene Representation (Shi et al., 2024; Qin et al., 2024; Zuo et al., 2024). Due to the high computational and memory demands of current language-embedded scene representations, Shi et al. (Shi et al., 2024) proposed a quantization scheme that augments 3D Gaussian with streamlined language embeddings instead of the original high-dimensional embeddings. This method also mitigated semantic ambiguity and enhanced the precision of open-vocabulary querying by smoothing out semantic features across different views, guided by uncertainty values. ii) Scene Understanding and Editing (Zhou et al., 2024b; Ye et al., 2024; Cen et al., 2023). Feature 3DGS (Zhou et al., 2024b) integrated 3D GS with feature field distillation from 2D foundation models. By learning a lower-dimensional feature field and applying a lightweight convolutional decoder for upsampling, Feature 3DGS achieved faster training and rendering speeds while enabling high-quality feature field distillation, supporting applications like semantic segmentation and language-guided editing. iii) Spatiotemporal Modeling (Yang et al., 2024b; Lin et al., 2024b) (task-centric summary and comparisons are given in Sec. 5.2). To capture the complex spatial and temporal dynamics of 3D scenes, Yang et al. (Yang et al., 2024b) conceptualized spacetime as a unified entity and approximates the spatiotemporal volume of dynamic scenes using a collection of 4D Gaussians. The proposed 4D Gaussian representation and corresponding rendering pipeline are capable of modeling arbitrary rotations in space and time and allow for end-to-end training.
@@ 4.6. Hybrid Representation
Next we showcase various fascinating uses of 3D GS with specially devised structured information. i) Facial Expression Modeling. Considering the challenge of creating high-fidelity 3D head avatars under sparse view conditions, Gaussian Head Avatar (Xu et al., 2024c) introduced controllable 3D Gaussians and an MLP-based deformation field. Concretely, it captured detailed facial expressions and dynamics by optimizing neutral 3D Gaussians alongside the deformation field, thus ensuring both detail fidelity and expression accuracy. ii) Spatiotemporal Modeling (task-centric summary and comparisons are given in Sec. 5.2). Yang et al. (Yang et al., 2024d) proposed to reconstruct dynamic scenes with deformable 3D Gaussians. The deformable 3D Gaussians are learned in a canonical space, coupled with a deformation field (i.e., a spatial MLP) that models the spatial-temporal dynamics. The proposed method also incorporated an annealing smoothing training mechanism to enhance temporal smoothness without additional computational costs. iii) Style Transfer. Saroha et al. (Saroha et al., 2024) proposed GS in style, an advanced approach for real-time neural scene stylization. To maintain a cohesive stylized appearance across multiple views without compromising on rendering speed, they used pre-trained 3D Gaussians coupled with a multi-resolution hash grid and a small MLP to produce stylized views. In a nutshell, incorporating structured information can serve as a complementary part for adapting to tasks that are incompatible with the sparsity and disorder of 3D Gaussians.